### v2

# InstructMPC: A Human-LLM-in-the-Loop Framework for Context-Aware Control

Abstract
Model Predictive Control (MPC) is a powerful control strategy widely utilized in domains like energy management, building control, and autonomous systems. However, its effectiveness in real-world settings is challenged by the need to incorporate context-specific predictions and expert instructions, which traditional MPC often neglects. We propose InstructMPC, a novel framework that addresses this gap by integrating real-time human instructions through a Large Language Model (LLM) to produce context-aware predictions for MPC. Our method employs a Language-to-Distribution (L2D) module to translate contextual information into predictive disturbance trajectories, which are then incorporated into the MPC optimization. Unlike existing context-aware and language-based MPC models, InstructMPC enables dynamic human-LLM interaction and fine-tunes the L2D module in a closed loop with theoretical performance guarantees, achieving a regret bound of $O(\sqrt{T\log T})$ for linear dynamics when optimized via advanced fine-tuning methods such as Direct Preference Optimization (DPO) using a tailored loss function.
1 Introduction
Model Predictive Control (MPC) is a versatile control strategy widely applied in domains such as energy management [1, 2], building control [3], and autonomous systems [4, 5, 6]. These applications leverage MPC’s ability to optimize system’s behavior over a predictive horizon, making it a cornerstone of modern control engineering.
In real-world scenarios, however, the predictions used in MPC are not solely derived from static models. Instead, they often depend on concrete, context-specific scenarios or are determined by experts to be near-optimal. For example, in building energy management of a university campus, MPC must anticipate and adapt to scenarios such as impending paper submission deadlines causing computational load surges, scheduled public events increasing demand, and forecasted weather changes affecting PV generation. This context dependency underscores a critical question: “How can we incorporate in-context instructions into MPC?” Key challenges in addressing this issue revolve around adapting to evolving human instructions and ensuring generalizability in changing environments. The MPC system must be capable of dynamically adjusting its behavior in response to shifting human inputs and contexts, which demands a high level of flexibility and responsiveness to maintain effectiveness. Furthermore, the learning-based integration of these instructions must be designed to adapt to time-varying conditions, allowing the system to not only respond to immediate changes but also to generalize its learning effectively, ensuring consistent performance over time. This combination of adaptability and robust generalization is essential for overcoming the complexities presented by dynamic human interactions and fluctuating environmental factors.
To tackle these challenges, we introduce InstructMPC, a novel framework that employs a human-LLM-in-the-loop approach to form a closed-loop system. InstructMPC takes in-context instructions from humans, processes them through a Large Language Model (LLM) to generate context-aware predictions, and feeds these into the downstream MPC module. During online implementation, the model loss from MPC is propagated back to the LLM, which is fine-tuned using techniques such as Proximal Policy Optimization (PPO) [7] and Direct Preference Optimization (DPO) [8] to enhance adaptability and generalizability across scenarios.
Recent advancements in MPC have explored contextual and language-based enhancements to improve performance in autonomous systems and personalized applications.
Context-aware MPC for autonomous systems
In autonomous systems, context-aware MPC leverages environmental and semantic information to enhance control performance. For instance, semantically informed MPC [9] employs deep neural networks to encode semantic labels (e.g., “find a couch”) into cost maps, guiding navigation in unknown environments with continuous control, achieving an improvement in success rates through mid-level visual cues. However, it focuses on discrete, predefined contexts without real-time human input. Similarly, [10] adapts MPC parameters to environmental contexts (e.g., rain) using contextual Bayesian optimization. While effective for model and objective tuning, it lacks mechanisms for incorporating real-time instructions from humans. Additionally, designed for crowded environments, the framework in [11] integrates contextual cues like human poses and activities into MPC. It prioritizes computational efficiency but does not generalize to in-context instructions beyond navigation tasks.
Language-Based MPC Personalization
Language-based approaches harness natural language to personalize MPC systems for improved usability and flexibility. For instance, the LanguageMPC in [12] utilizes LLMs as decision-makers in autonomous driving, combining LLM reasoning with low-level controllers to enhance safety, efficiency, and generalizability in complex scenarios, though it does not support real-time human interaction. Meanwhile, the ChatMPC [13] processes natural language via BERT to update MPC specifications dynamically, reducing the data burden for personalization; however, it lacks state-of-the-art human-LLM interaction and theoretical performance guarantees.
The InstructMPC framework integrates real-time human instructions through a closed-loop human-LLM system for dynamic adaptation and provides a theoretical performance guarantee for linear dynamics using DPO, ensuring features absent in existing context-aware and language-based approaches. Our contributions are two-fold:
  1. 1.
    InstructMPC Framework: InstructMPC dynamically tunes MPC by integrating real-time, human-provided in-context instructions through a Language-to-Distribution (L2D) module (e.g., LLMs) and fine-tuning the L2D module in a closed loop. The novelty lies in enabling the MPC controller to leverage contextual information often ignored in traditional control models. By transforming high-level instructive information into predicted disturbance trajectories, InstructMPC enhances the adaptability of classic MPC.
  2. 2.
    Theoretical Guarantees: For linear dynamics, we provide theoretical performance (Theorem 4.1) guarantees for closed-loop LLM fine-tuning using Direct Preference Optimization (DPO). Unlike traditional online optimization methods that directly minimize regret, our approach decouples the DPO loss from the regret objective, which relies on MPC cost information and is challenging to compute online due to unknown future disturbances. In particular, Theorem 4.1 suggests a tailored loss function for LLM fine-tuning, achieving a regret bound of $O(\sqrt{T\log T})$ as derived in Corollary 4.1.
2 Preliminaries and Problem Formulation
Denote $[T]\coloneqq\{0,1,\ldots,T-1\}$. We consider a finite horizon linear dynamical system
$x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\quad t\in[T]$ (1)
where $x_{t}\in\mathbb{R}^{n},u_{t}\in\mathbb{R}^{m}$ denote the system state and control input at each time $t\in[T]$; $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are system matrices. The disturbance $w_{t}\in\mathbb{R}^{n}$ is unknown to the controller at time $t\in[T]$. The controller’s goal is to minimize the following quadratic costs:
$J^{\star}\coloneqq\min_{(u_{t}:t\in[T])}$ $\sum_{t=0}^{T-1}(x_{t}^{\top}Qx_{t}+u_{t}^{\top}Ru_{t})+x_{T}^{\top}Px_{T},$ (2a)
[…]
where $Q,R\succ 0$ are positive definite matrices, and $P$ is the solution to the following discrete algebraic Riccati equation (DARE), [^1] To simplify the presentation of our main results, we fix the terminal cost in (2a) to be $P$. The arguments extend to more general terminal costs as well, since the overall cost only differs by an $O(1)$ term.
$P=Q+A^{\top}PA-A^{\top}PB(R+B^{\top}PB)^{-1}B^{\top}PA.$ (3)
The disturbance $w_{t}$ is bounded by a constant $W>0$ (i.e. $w_{t}\in\mathcal{W}\coloneqq\{w:\|w\|\leq W\}$) for all $t\in[T]$. Futhermore, the system $(A,B)$ is stabilizable [14].
Assumption 1.
There exists $K\in\mathbb{R}^{m\times n}$ such that $A-BK$ has a spectral radius $\rho$ less than $1$. [^2] Thus, the Gelfand’s formula implies that there exist $C>0$, $\rho\in(0,1)$ such that $\|A-BK\|^{t}\leq C\rho^{t}$ for all $t\geq 0$.
2.1 Model Predictive Control
Fix a prediction horizon to be an integer $k$ and define $\mathcal{T}\coloneqq\min\{t+k-1,T-1\}$. We consider the setting in which, at each time $t\in[T]$, $k$-step predictions $\hat{w}_{t:\mathcal{T}|t}\coloneqq(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})\in\mathcal{W}^{k}$ corresponding to future disturbances $(w_{t},\ldots,w_{\mathcal{T}})$ are available.
In this context, model predictive control naturally incorporates these predictions into its optimization framework. At each time $t\in[T]$, the controller solves the following optimization problem given the current state $x_{t}$ and disturbance predictions in $\hat{w}_{t:\mathcal{T}|t}$:
$u_{t:\mathcal{T}}^{\textsc{MPC}}$ $\coloneqq\arg\min_{u}\Big{(}\sum_{\tau=t}^{\mathcal{T}}(x_{\tau}^{\top}Qx_{\tau}+u_{\tau}^{\top}Ru_{\tau})+x^{\top}_{\mathcal{T}+1}Px_{\mathcal{T}+1}\Big{)}$
[…]
Only the first element of $u_{t:\mathcal{T}}^{\textsc{MPC}}$ is applied as a control input to the system, and the rests are discarded. However, in many real-world scenarios, the disturbances are hard to predict or measure [15, 16]. In this paper, we propose a novel framework that leverages contextual information to predict future disturbances and incorporates these predictions into the decision-making process.
2.2 Contextual MPC with Model Fine-Tuning
[…]
Assumption 2.
The NN model $g_{\theta}$ is differentiable and affine in $\theta$ and its gradient satisfies $\|\nabla_{\theta}g_{\theta}\|\leq L$ for some constant $L>0$.
A hyper-parameter $\theta_{t}$ is updated at $t\in[T]$. We evaluate performance via the following regret, casting the problem as an online optimization over $\theta_{1:T}\coloneqq(\theta_{t}:t\in[T])$:
$J(\theta_{1:T})-\min_{\theta\in\Theta}J(\theta),$ (5)
where $J(\theta_{1},\ldots,\theta_{T})$ and $J(\theta^{\star})$ are the quadratic costs defined in (2a) induced by the MPC actions in (2.1) corresponding to the hyper-parameters $\theta_{1},\ldots,\theta_{T}$ and the optimal hyper-parameter $\theta^{\star}$, respectively. In other words, the regret in Equation 5 measures the difference between the cumulative loss incurred by $(\theta_{t}:t\in[T])$, and the cumulative loss incurred by the optimal $\theta^{\star}$ in hindsight. Next, we introduce the InstructMPC framework for contextual MPC.
3 InstructMPC
In this section, we introduce InstructMPC, a framework that enables MPC controller to leverage external contextual information to effectively manage environmental disturbances. Fig. 1 presents the overall architecture of InstructMPC. In the following subsections, we explain the core component of InstructMPC, Language to Distribution (L2D) module in details.
Figure 1: System framework of InstructMPC. The blue lines represent interactions between InstructMPC and the environment where InstructMPC receives the state $x_{t}$, and outputs the control input $u_{t}$. The black lines represents the information loop, within which external contextual information is passed to the L2D module $g_{\theta}$ to produce predicted disturbances $\hat{w}_{t:\mathcal{T}|t}$. Then, the MPC controller $\pi_{\mathrm{MPC}}$ utilizes $\hat{w}_{t:\mathcal{T}|t}$ and the current state $x_{t}$ to determine a control input $u_{t}^{\mathrm{MPC}}$ via (2.1). After executing the MPC control input $u_{t}^{\mathrm{MPC}}$, the environment reveals the true disturbance $w_{t}$. The discrepancy between $w_{t}$ and $\hat{w}_{t}$ is then sent back to the L2D module as a loss signal.
3.1 Language to Distribution (L2D) Module
During the execution of InstructMPC, the L2D module $g_{\theta}$ predicts future environmental disturbances $\hat{w}_{t:\mathcal{T}|t}=(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})$ based on contextual information. After each control step, the module refines its parameters $\theta$ by comparing these predictions with the actual revealed disturbances.
At each time $t$, given contextual information $c_{t}\in\mathcal{C}$, the module first assigns a probability $p(s|c_{t})$ to each scenario $s\in\mathcal{S}$. For each scenario $s$, the module retrieves a disturbance trajectory $w^{s}_{t:\mathcal{T}}=(w^{s}_{t|t},\ldots,w^{s}_{\mathcal{T}|t})$ based on historic data. The predicted disturbance trajectory is computed by taking a weighted combination of trajectories $(w^{s}_{t:\mathcal{T}}:s\in\mathcal{S})$:
$\hat{w}_{t:\mathcal{T}|t}=\sum_{s\in\Omega}p(s|c_{t})w^{s}_{t:\mathcal{T}}.$
3.2 InstructMPC Framework
for $t=0,\dots,T-1$ do
       $\mathcal{T}\leftarrow\min(t+k-1,T-1)$
       Get predictions from $g_{\theta}$ with instruction $c_{t}$: $(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})\leftarrow g_{\theta_{t}}(c_{t})$ Generate an action $u_{t}$ by solving MPC in (2.1) Update
$x_{t+1}$ $=Ax_{t}+Bu_{t}+w_{t}$
$\theta_{t+1}$ $=\theta_{t}-\eta_{t}\nabla_{\theta}L_{t-k+1}(\theta_{t-k+1})$
end for
Algorithm 1 InstructMPC
The L2D module is a neural network $g_{\theta}:\mathcal{C}\to\mathcal{W}^{k}$, which is updated at every time $t$ through a loss function $L_{t-k+1}:\Theta\times\mathcal{W}^{2k}\rightarrow\mathbb{R}$, which also depends on disturbances $w_{t:\mathcal{T}}$ and predictions $\hat{w}_{t:\mathcal{T}|t}$. An example of such a loss function can be found in Corollary 4.1 in Section 4. At every time $t\in[T]$, the parameter $\theta$ of L2D module is updated according to:
$\theta_{t+1}=\theta_{t}-\eta_{t}\nabla_{\theta}L_{t-k+1}(\theta_{t-k+1}),$ (6)
where $\eta_{t}\nabla_{\theta}L_{t-k+1}$ denotes the gradient of $L_{t-k+1}$ with respect to $\theta$ and $\eta_{t}$ is a time-varying learning rate. The gradient descent rule above is widely used to fine-tune NN models, such as Direct Preference Optimization (DPO) [8], Proximal Policy Optimization (PPO) [7], and Supervised Fine-Tuning (SFT) [19]. The overall procedure of InstructMPC is summarized in Algorithm 1. Since $\nabla_{\theta}L_{t}(\theta_{t})$ is available only after $(w_{t},\ldots,w_{\mathcal{T}})$ are revealed at time $t+k$, a delayed gradient is used in the update rule (6). For convenience, we define $L_{t-k+1}=0$ for $t<k$. In practice, when $t<k$, the NN model does not update due to insufficient observed data.
4 Theoretical Guarantee
Our main result in this section explores the online optimization of $(\theta_{t}:t\in[T])$ in the online fine-tuning process. We make the following standard assumption on the loss function $L_{t}$.
Assumption 3.
The gradient $\nabla_{\theta}L_{t}(\theta)$ is uniformly bounded, i.e., there exists $G>0$ such that by $\|\nabla_{\theta}L_{t}(\theta)\|\leq G$ for all $\theta\in\Theta$.
Note that the loss function used for updating $g_{\theta}$ may not be aligned with regret-optimal objective that requires the knowledge of future ground-truth predictions $(w_{t}:t\in[T])$ in (2.1). To quantify this mismatch, we define the loss discrepancy below.
Definition 1.
The loss discrepancy between two loss functions $L_{1}(\theta)$ and $L_{2}(\theta)$, $\textsf{LD}(L_{1},L_{2})$ is defined as
$\textsf{LD}(L_{1},L_{2})\coloneqq\sup_{\theta\in\Theta}\left\|\nabla_{\theta}L_{1}(\theta)-\nabla_{\theta}L_{2}(\theta)\right\|.$
The theorem below provides a regret bound (see Equation 5) for the proposed InstructMPC, which interacts with the dynamic system described via Equation 1.
Theorem 4.1.
Under Assumption 2,3, if the learning rate $\eta_{t}$ is non-increasing, then
$J(\theta_{1:T})-$ $J(\theta^{\star})\leq\frac{D^{2}}{2\eta_{T-1}}+\left(k-\frac{1}{2}\right)G^{2}\sum_{t=0}^{T-1}\eta_{t}$
$+D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$ (7)
Furthermore, if we choose $\eta_{t}=\frac{D}{G\sqrt{2(2k-1)(t+1)}}$,
$J(\theta_{1:T})-$ $J(\theta^{\star})\leq 2GD\sqrt{(k-\frac{1}{2})T}$
$+D\sum_{t=0}^{T-1}\textsf{LD}\left(L_{t},\psi_{t}^{\top}H\psi_{t}\right)+(k-1)GD,$
where we define
$\psi_{t}(\theta)\coloneqq\sum_{\tau=t}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pg_{\theta}^{(\tau-t+1)}(c_{t}).$
Proof.
By Theorem 3.2 in [20], given predictions $\hat{w}_{t:\mathcal{T}|t}$, the MPC solution $u_{t}^{\textsc{MPC}}$ to the problem defined in (2.1) is
$u_{t}^{\textsc{MPC}}=-(R+B^{\top}PB)^{-1}B^{\top}\Big{(}PAx_{t}+\sum_{\tau=t}^{\mathcal{T}}(F^{\top})^{\tau-t}P\hat{w}_{\tau|t}\Big{)},$
where $F\coloneqq A-B(R+B^{\top}PB)^{-1}B^{\top}PA\coloneqq A-BK$. Thus, applying Lemma 1 twice, we obtain
$J(\theta_{1:T})-J(\theta^{\star})=\sum_{t=0}^{T-1}\psi_{t}(\theta_{t})^{\top}H\psi_{t}(\theta_{t})-\psi_{t}(\theta^{\star})^{\top}H\psi_{t}(\theta^{\star}),$
where $\psi_{t}(\theta)$ is defined in Theorem 4.1. For notational simplicity, we define $\psi_{t}\coloneqq\psi_{t}(\theta_{t})$ and $\phi_{t}\coloneqq\psi_{t}(\theta^{\star})$. By our model assumption, $Q,R\succ 0$ implies $H\succeq 0$, using the convexity of $x^{\top}Hx$,
$\frac{1}{2}\sum_{t=0}^{T-1}\psi_{t}^{\top}H\psi_{t}-\phi_{t}^{\top}H\phi_{t}\leq\sum_{t=0}^{T-1}\psi_{t}^{\top}H(\psi_{t}-\phi_{t})=$
$\sum_{t=0}^{T-1}\psi_{t}^{\top}H\left(\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}P\left(g_{\theta^{\star}}^{(\tau-t+1)}(c_{t})-g_{\theta_{t}}^{(\tau-t+1)}(c_{t})\right)\right).$
Applying Assumption 2, since $g_{\theta}$ is affine in $\theta$, continuing from above,
$J(\theta_{1:T})-J(\theta^{\star})\leq$
$-2\sum_{t=0}^{T-1}\psi_{t}^{\top}H\left(\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}P\nabla_{\theta}g_{\theta_{t}}^{(\tau-t+1)}(c_{t})^{\top}\right)(\theta_{t}-\theta^{\star})$
$=\sum_{t=0}^{T-1}\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})^{\top}(\theta_{t}-\theta^{\star}),$ (8)
where we denote $\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})\coloneqq\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})|_{\theta=\theta_{t}}$, and have used Assumption 2. Denote the gradient of the loss function as $\nabla_{\theta}L_{t-k+1}(\theta_{t-k+1})$ as $l_{t-k+1}$. It follows that the RHS of (8) can be rewritten as
$\sum_{t=0}^{T-1}\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})^{\top}(\theta_{t}-\theta^{\star})$
$=$ $\sum_{t=0}^{T-1}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})+\sum_{t=0}^{T-1}(l_{t}-l_{t-k+1})^{\top}(\theta_{t}-\theta^{\star})$
$\quad+\sum_{t=0}^{T-1}(\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})^{\top}-l_{t}^{\top})(\theta_{t}-\theta^{\star}).$ (9)
The bound in (7) is obtained by bounding the terms in Equation 9 separately, and the details are relegated to Appendix A. ∎
Furthermore, for a wide range of loss functions $(L_{t}:t\in[T])$, as examplified in Corollary 4.1 below, they exhibit a bounded discrepancy from the MPC cost with the discrepancy decaying exponentially as $k$ increases, i.e., $\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k}$.
Corollary 4.1.
Under Assumption 1, for the case when $L_{t}(\theta)=\hat{\psi}_{t}(\theta)^{\top}H\hat{\psi}_{t}(\theta)$ with
$\hat{\psi}_{t}(\theta)\coloneqq\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pg_{\theta}^{(\tau-t+1)}(c_{t}),$
there exist constants $C$ and $10<\rho<$ 1 such that for all $t\in[T]$,
$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k},$
where $\psi_{t}(\theta)$ is defined in Theorem 4.1.
For the particular choice of $L_{t}$ in Corollary 4.1, the term $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})$ in the bound (7) of Theorem 4.1 simplifies to $CDT\rho^{k}$. This provides an insightful guideline for selecting $k$. Specifically, by setting $k=\log_{1/\rho}T,$ we can achieve a regret bound of $O(\sqrt{T\log T})$.
5 Numerical Experiment
In this section, we present a detailed demonstration of how the L2D module is fine-tuned in Subsection 5.1. We implement the L2D module with LLaMa-8B [21], a causal language model. Causal language models predict the next token in a sequence and return prediction scores for each vocabulary token, which aligns well with L2D’s requirement of generating scenario probabilities. To obtain these probabilities, we use the following prompt template: “[task description]. Given the current context: [context description], the expected scenario is [MASK]." For a predefined set of scenarios ($\mathcal{S}$ in Subsection 3.1), we evaluate the prediction score assigned by the model by inserting each candidate token into the [MASK] position. These scores are then normalized using a softmax function to produce a probability distribution.
In Subsection 5.2 and Subsection 5.3, we demonstrate the effectiveness of InstructMPC framework in two scenarios. The first is a robot tracking task, where an agent is required to follow an unknown trajectory. In this setting, the target position is revealed only immediately before the control decision is made. Moreover, environmental disturbances, such as wind force, are described textually and passed through the L2D module to assist the MPC controller. We plot the resulting trajectory to directly demonstrate the impact of contextual information. The second scenario is a real-world energy management task, in which human-provided instructions, such as news and scheduled appointments, are used to predict external disturbances. In this case, we further extend InstructMPC to operate on a nonlinear dynamical system, we also adopt a piecewise linear cost instead of the quadratic cost defined in eq. 2a, demonstrating the generalizability of InstructMPC in diverse system dynamics and different cost functions.
5.1 Fine-Tuning L2D Module
Fine-Tuning L2D Module. We first define the preferences among scenarios based on observed disturbances. Given the the prediction horizon $k$, and $\mathcal{T}\coloneqq\min(t+k-1,T-1)$, after each $k$ control steps, the true disturbance sequence $w_{t:\mathcal{T}}$ is revealed. For each scenario $s\in\mathcal{S}$, we measure the distance between $w_{t:\mathcal{T}}$ and $w_{t:\mathcal{T}}^{s}$ as:
$\mathrm{dist}(w_{t:\mathcal{T}},w_{t:\mathcal{T}}^{s})\coloneqq\|w_{t:\mathcal{T}}-w_{t:\mathcal{T}}^{s}\|.$
Given the context $c_{t}$ at time $t$, we say scenario $s_{1}$ is more likely to occur than scenario $s_{2}$, denoted as $s_{1}\succ s_{2}$, if $\mathrm{dist}(w_{t:\mathcal{T}},w_{t:\mathcal{T}}^{s_{1}})<\mathrm{dist}(w_{t:\mathcal{T}},w_{t:\mathcal{T}}^{s_{2}})$. This yields a dataset of pairwise comparisons:
$\mathcal{D}=\{(c_{t},s_{w,t},s_{l,t})|t\in\mathcal{T}_{\text{obs}};s_{w},s_{l}\in\mathcal{S};s_{i}\neq s_{j}\},$
where $\mathcal{T}_{\text{obs}}\in\{1,\ldots,T\}$ denotes the subset of time step at which observations and comparisons are collected, $s_{w,t}$ and $s_{l,t}$ denote the preferred and dispreferred scenario at time $t$ respectively. In practice, the parameter adjustment loop operates at a different timescale than the control loop, we tune the L2D module in batches. During both the robot tracking and energy management experiments, the L2D module is fine-tuned using DPO [8] once the size of pairwise comparison dataset $\mathcal{D}$ reaches a predefined threshold. Specifically, the parameter $\theta$ is updated when $|\mathcal{D}|=240$ in the robot tracking task and $|\mathcal{D}|=500$ in the energy management scenario.
5.2 Application 1: Robot Tracking
Problem description. We examine a 2D robotic tracking scenario [22, 20] where a controller follows an unknown, fixed astroid path, which is
$y_{t}\coloneqq\begin{bmatrix}2\sin^{3}(t/38.2)\\ 2\cos^{3}(t/38.2)\end{bmatrix},\quad t\in[T].$
Let $p_{t}\in\mathbb{R}^{2}$ and $v_{t}\in\mathbb{R}^{2}$ denote the location and the velocity of robot at time $t$. $p_{t+1}$ depends on previous location $p_{t}$ and velocity $v_{t}$ such that $p_{t+1}=p_{t}+0.2v_{t}$. The velocity is influenced by both the control input $u_{t}$ and the external disturbances, specifically the wind force represented by $Z_{t}$. Let $x_{t}\coloneqq q_{t}-y_{t}$ represent the deviation between the robot’s trajectory and the target trajectory. The system dynamics can be represented by:
$\begin{bmatrix}x_{t+1}\\ v_{t+1}\end{bmatrix}=A\begin{bmatrix}x_{t}\\ v_{t}\end{bmatrix}+Bu_{t}+w_{t},$
where
$A\coloneqq\begin{bmatrix}1&0&0.2&0\\ 0&1&0&0.2\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},B\coloneqq\begin{bmatrix}0&0\\ 0&0\\ 0.2&0\\ 0&0.2\end{bmatrix},$
$w_{t}\coloneqq Ay_{t}-y_{t+1}+Z_{t},$
$Z_{t}\coloneqq\begin{bmatrix}0\!\!\!&0\!\!\!&-0.2Z_{t}^{(1)}\!\!\!&-0.2Z_{t}^{(2)}\end{bmatrix}^{\top}\!\!\!,$
where $Z_{t}$ is a random variable representing the random wind force. Specifically, during the experiment, we randomly select a subset of $[T]$, denote as $T_{\text{sub}}=\{20,25,32,35,40,84,133,145,158,215\}$, and let
$\begin{cases}Z^{(1)}_{t},Z^{(2)}_{t}\sim\mathcal{U}(-45,45)&t\in T_{\text{sub}}\\ Z^{(1)}_{t},Z^{(2)}_{t}\sim\mathcal{U}(-2,2)&\text{otherwise}\end{cases}$
To track the trajectory, the controller uses quadratic cost in Equation 2a with
$Q\coloneqq\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},\ \ R\coloneqq\begin{bmatrix}10^{-2}&0\\ 0&10^{-2}\end{bmatrix}.$
Experimental results. In our first experiment, we plot the robot trajectory to demonstrate the impact of external contextual information, in Fig. 2, a zoom-in figure is used to demonstrate the exact impact of instructions, We compare three settings:
  1. 1.
    Classic MPC: An MPC baseline using (2.1) where disturbance predictions are set to zero ($\hat{w}_{t:\mathcal{T}|t}=0$) for all $t$, equivalent to a Linear Quadratic Regulator (LQR) for the problem (2a)-(2b).
  2. 2.
    Zero-Shot Prompting: MPC with predictions generated by a zero-shot prompted LLaMA-8B [21] model.
  3. 3.
    InstructMPC (with fine-tuning): Our proposed method using LLaMA-8B [21] fine-tuned in a closed-loop with DPO [8].
We observe that the baseline (classic MPC without L2D predictions) tends to deviate from the trajectory due to its inability to predict disturbances from external information. In contrast, both methods with L2D predictions improve the performance of the robot, with the fine-tuned version demonstrating a closer robot path to the designed trajectory.
Figure 2: Application 1: Robot Tracking (Section 5.2). Average robot tracking trajectories of three methods: 1. MPC in (2.1) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2. MPC with predictions from a zero-shot prompting LLaMA-8B [21] model; 3. InstructMPC with LLaMA-8B [21] fine-tuned in a closed-loop. For each method, we ran $20$ episodes, and plot the average robot trajectory. The shadow region denotes the variance of trajectories.
5.3 Application 2: Building Energy Management
Problem description. We consider the problem of energy management for a campus building. The system consists of a battery for energy storage and a series of photovoltaic (PV) arrays to harvest the sunlight. Let $x_{t}\in[0,1]$ denote the state of charge (SoC) of the battery at time $t$, and $u_{t}\in[0,1]$ represent the charging rate, $u_{t}=1$ corresponds to charging at the maximum allowable power, and $u_{t}=0$ indicates no charging. The nonlinear system dynamics can be represented by: $x_{t+1}=f(x_{t},u_{t})+w_{t},$ where $f:\mathbb{R}^{2}\to\mathbb{R}$ captures the system’s inherent battery dynamics, $w_{t}\in\mathbb{R}$ represents the external disturbance, which can be attributed to the variations in solar irradiance or fluctuations in electricity demand. These disturbances are inferred from contextual information, including weather forecasts, human activity schedules, and natural language instructions. Examples of such contextual information include: “The CDC conference deadline is approaching, and many people are running experiments," or “I will be fine-tuning my LLaMA-1B model at 3 PM today, which will take approximately two hours." The objective of the energy management system is to minimize the total cost of purchasing electricity from the grid. This cost is computed using a piecewise linear function that reflects real-world peak and valley electricity prices.
Figure 3: Application 2: Building Energy Management (Section 5.3). Daily electricity cost comparison of three approaches: 1. MPC in (2.1) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2. MPC with predictions from a zero-shot prompting LLaMA-8B [21] model; 3. InstructMPC with LLaMA-8B [21] fine-tuned in a closed-loop.
Experimental results. We show the performance of InstructMPC in an energy management scenario using a synthetic workload and real-world weather data. Each energy-consuming task arrives at a constant rate of $1$ task per unit time, and its energy demand is drawn from a uniform distribution $\mathcal{U}(1,5)$, with each unit corresponding to $0.5$ kWh. For PV generation, we randomly sample $500$ days of real weather data and model the corresponding PV generation using PVLIB simulator [23], and simulate three different MPC settings described above.
The results are presented in Fig. 3. The left part shows the daily electricity costs over $500$ days for all three methods, demonstrating the effectiveness of incorporating contextual predictions and model parameter adjustment. The upper-right part provides a summary table comparing the total electricity cost and the corresponding percentage reduction, while the bottom-right part provides a closer view of daily costs over $40$ days, further illustrating the cost savings by InstructMPC.
6 Conclusion
We have presented the InstructMPC framework that integrates real-time human instructions into Model Predictive Control (MPC) via a human-LLM-in-the-loop approach. This work demonstrates how leveraging contextual information can significantly improve MPC’s predictive accuracy and performance in complex, real-world applications. By dynamically generating context-aware disturbance predictions and refining them using advanced techniques like Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO), InstructMPC offers improved adaptability and performance across a wide range of applications. For linear dynamics, we established a robust performance guarantee, proving that IMPC achieves a regret bound of $O(\sqrt{T\log T})$. In future work, we aim to extend these results to nonlinear systems and further explore the integration of richer contextual cues to enhance adaptability and robustness in complex environments.
Appendix A Proof of Theorem 4.1
The following lemma characterizing the quadratic cost gap is useful.
Lemma 1 (Lemma 13 in [24]).
For any $\psi_{t}\in\mathbb{R}^{n}$, if at each time $t\in[T]$, a controller $\pi$ implements a control input
$u_{t}^{\pi}=-(R+B^{\top}PB)^{-1}B^{\top}\Big{(}PAx_{t}+\sum_{\tau=t}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\psi_{t}\Big{)},$
then the difference between the optimal cost $J^{\star}$ and the algorithm cost $J(\pi)$ is given by:
$J(\pi)-J^{\star}=\sum_{t=0}^{T-1}\psi_{t}^{\top}H\psi_{t},$
where $H\coloneqq B(R+B^{\top}PB)^{-1}B^{\top}$ and $F\coloneqq A-HPA$.
Now, we bound the terms in Equation 9 separately:
$\sum_{t=0}^{T-1}\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})^{\top}(\theta_{t}-\theta^{\star})$
$=$ $\underbracket{\sum_{t=0}^{T-1}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})}_{\text{(a)}}+\underbracket{\sum_{t=0}^{T-1}(l_{t}-l_{t-k+1})^{\top}(\theta_{t}-\theta^{\star})}_{\text{(b)}}$
$\quad+\sum_{t=0}^{T-1}(\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})^{\top}-l_{t}^{\top})(\theta_{t}-\theta^{\star}).$
First, based on the gradient update rule in (6) of InstructMPC,
$\|\theta_{t+1}-\theta^{\star}\|^{2}$
$=$ $\|\theta_{t}-\eta_{t}l_{t-k+1}-\theta^{\star}\|^{2}$
$=$ $\|\theta_{t}-\theta^{\star}\|^{2}-2\eta_{t}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})+\eta_{t}^{2}\|l_{t-k+1}\|^{2}.$ (10)
Then, rearranging the terms in Equation 10 and noting that the gradients are bounded by Assumption 3,
$2l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})\leq\frac{\|\theta_{t}-\theta^{\star}\|^{2}-\|\theta_{t+1}-\theta^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (11)
Summing above from $t=0$ to $T-1$, for (a),
22 22 $\sum_{t=0}^{T-1}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})$
$\leq$ $\sum_{t=0}^{T-1}\frac{\|\theta_{t}-\theta^{\star}\|^{2}-\|\theta_{t+1}-\theta^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}$
$\leq$ $\sum_{t=0}^{T-1}\|\theta_{t}-\theta^{\star}\|^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=0}^{T-1}\eta_{t}$
$\leq$ $\frac{D^{2}}{\eta_{T-1}}+G^{2}\sum_{t=0}^{T-1}\eta_{t}.$ (12)
Furthermore, the second term in Equation 9 can be bounded from above as
[…]
$=$ $\sum_{t=0}^{T-k}l_{t}^{\top}(\theta_{t}-\theta_{t+k-1})+\sum_{t=T-k+1}^{T-1}l_{t}^{\top}(\theta_{t}-\theta^{\star})$
$\leq$ $\sum_{t=0}^{T-k}(G\|\theta_{t}-\theta_{t+k-1}\|)+(k-1)GD$
$=$ $(k-1)GD+G\sum_{t=0}^{T-k}\left\|\sum_{\tau=t}^{t+k-2}\eta_{\tau}l_{\tau-k+1}\right\|$
$\leq$ $(k-1)GD+G^{2}\sum_{t=0}^{T-k}\sum_{\tau=t}^{t+k-2}\eta_{\tau}$
$\leq$ $(k-1)GD+(k-1)G^{2}\sum_{t=0}^{T-1}\eta_{t}.$ (13)
Recalling (8),
$J(\theta_{1:T})-J(\theta^{\star})$
$\leq$ $\frac{D^{2}}{2\eta_{T-1}}+\frac{G^{2}}{2}\sum_{t=0}^{T-1}\eta_{t}+(k-1)GD+(k-1)G^{2}\sum_{t=0}^{T-1}\eta_{t}$
$+$ $\sum_{t=0}^{T-1}\|\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})-l_{t}\|\|\theta_{t}-\theta^{\star}\|$ (14)
$\leq$ $\frac{D^{2}}{2\eta_{T-1}}+(k-\frac{1}{2})G^{2}\sum_{t=0}^{T-1}\eta_{t}+(k-1)GD$
$+$ $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t}),$ (15)
where we have used (12) and (13) to derive  (14); the last inequality (15) holds by the loss discrepancy in Definition 1. Finally, applying $\eta_{t}={D}/\left({G\sqrt{2(2k-1)(t+1)}}\right)$ to (15),
$J(\theta_{1:T})-J(\theta^{\star})$
$=$ $GD\sqrt{(k-\frac{1}{2})T}+\sqrt{k-\frac{1}{2}}\frac{GD}{2}\sum_{t=0}^{T-1}\frac{1}{\sqrt{t+1}}$
$+$ $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD$
$\leq$ $2GD\sqrt{(k-\frac{1}{2})T}+D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$ (16)
Inequality (16) in the second statement of Theorem 4.1 is obtained by using $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$.
Appendix B Proof of Corollary 4.1
Suppose $L_{t}(\theta)=\hat{\psi}_{t}(\theta)^{\top}H\hat{\psi}_{t}(\theta)$ where
$\hat{\psi}_{t}(\theta)\coloneqq\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pg_{\theta}^{(\tau-t+1)}(c_{t}).$
By Definition 1,
$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})$ $=\left\|\frac{\partial(\psi_{t}^{\top}H\psi_{t})}{\partial\theta}-\frac{\partial(\hat{\psi}_{t}^{\top}H\hat{\psi}_{t})}{\partial\theta}\right\|$
$=\left\|2(\psi_{t}-\hat{\psi}_{t})^{\top}H\frac{\partial\psi_{t}}{\partial\theta}\right\|$ (17)
$\leq 2\|H\|\left\|\psi_{t}-\hat{\psi}_{t}\right\|\left\|\frac{\partial\psi_{t}}{\partial\theta}\right\|.$ (18)
We have use the fact that $\frac{\partial\psi_{t}}{\partial\theta}=\frac{\partial\hat{\psi}_{t}}{\partial\theta}$ to get (17). Note that if $\mathcal{T}=T-1$, then $\psi_{t}=\hat{\psi}_{t}$, $\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})=0$. Thus, it suffices to restrict our analysis to the case when $\mathcal{T}=t+k-1$. It follows that
$\left\|\psi_{t}-\hat{\psi}_{t}\right\|$ $=\left\|\sum_{\tau=t+k}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}\right\|$
$\leq W\|P\|\sum_{\tau=t+k}^{T-1}\|F\|^{\tau-t}$ (19)
$\leq W\frac{\|P\|\|F\|^{k}}{1-\|F\|},$ (20)
where (19) follows from the Assumption 1 so that $\|w_{t}\|\leq W$ for all $t\in[T]$. Moreover, by Assumption 2, $\|\nabla_{\theta}g_{\theta_{t}}^{(\tau-t+1)}(c_{t})\|\leq L$,
$\left\|\frac{\partial\psi_{t}}{\partial\theta}\right\|$ $=\left\|\sum_{\tau=t}^{t+k-1}\left(F^{\top}\right)^{\tau-t}P\nabla_{\theta}g_{\theta_{t}}^{(\tau-t+1)}(c_{t})\right\|$
$\leq L\|P\|\sum_{\tau=t}^{t+k-1}\|F\|^{\tau-t}$
$\leq L\frac{\|P\|}{1-\|F\|}.$ (21)
Finally, combining (18),(20), and (21) we have
$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq 2\frac{LW\|P\|^{2}\|H\|\|F\|^{k}}{(1-\|F\|)^{2}}.$ (22)

### v3

@@
 Unlike existing context-aware and language-based MPC models, InstructMPC enables dynamic human-LLM interaction and fine-tunes the L2D module in a closed loop with theoretical performance guarantees, achieving a regret bound of $O(\sqrt{T\log T})$ for linear dynamics when optimized via advanced fine-tuning methods such as Direct Preference Optimization (DPO) using a tailored loss function.
-1 Introduction
-Model Predictive Control (MPC) is a versatile control strategy widely applied in domains such as energy management [1, 2], building control [3], and autonomous systems [4, 5, 6].
+I Introduction
+Model Predictive Control (MPC) is a versatile control strategy widely applied in domains such as energy management [1, 2, 3, 4], building control [5], and autonomous driving [6, 7].
 These applications leverage MPC’s ability to optimize system’s behavior over a predictive horizon, making it a cornerstone of modern control engineering.
@@
 For instance, the LanguageMPC in [13] utilizes LLMs as decision-makers in autonomous driving, combining LLM reasoning with low-level controllers to enhance safety, efficiency, and generalizability in complex scenarios, though it does not support real-time human interaction.
-Meanwhile, the ChatMPC [13] processes natural language via BERT to update MPC specifications dynamically, reducing the data burden for personalization; however, it lacks state-of-the-art human-LLM interaction and theoretical performance guarantees.
+Meanwhile, ChatMPC [14] uses BERT to classify natural language into a set of predefined circumstances, each linked to a predefined modification of MPC parameters; however, it is constrained by its predefined updating rule, and also lacks state-of-the-art human-LLM interaction and theoretical performance guarantees.
 The InstructMPC framework integrates real-time human instructions through a closed-loop human-LLM system for dynamic adaptation and provides a theoretical performance guarantee for linear dynamics using DPO, ensuring features absent in existing context-aware and language-based approaches.
@@
   2. 2.
-    Theoretical Guarantees: For linear dynamics, we provide theoretical performance (Theorem 4.1) guarantees for closed-loop LLM fine-tuning using Direct Preference Optimization (DPO).
+    Theoretical Guarantees: For linear dynamics, we provide theoretical performance (Theorem IV.1) guarantees for closed-loop LLM fine-tuning using Direct Preference Optimization (DPO).
 Unlike traditional online optimization methods that directly minimize regret, our approach decouples the DPO loss from the regret objective, which relies on MPC cost information and is challenging to compute online due to unknown future disturbances.
-In particular, Theorem 4.1 suggests a tailored loss function for LLM fine-tuning, achieving a regret bound of $O(\sqrt{T\log T})$ as derived in Corollary 4.1.
-2 Preliminaries and Problem Formulation
-Denote $[T]\coloneqq\{0,1,\ldots,T-1\}$. We consider a finite horizon linear dynamical system
-$x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\quad t\in[T]$ (1)
+In particular, Theorem IV.1 suggests a tailored loss function for LLM fine-tuning, achieving a regret bound of $O(\sqrt{T\log T})$ as derived in Corollary IV.1.
+II Preliminaries and Problem Formulation
+Denote $T]\coloneqq\{0,1,\ldots,T-1\}$. We consider a finite horizon linear dynamical system
+$x_{t+1}=Ax_{t}+Bu_{t}+w_{t},\quad t\in[T]\$(1)
 where $x_{t}\in\mathbb{R}^{n},u_{t}\in\mathbb{R}^{m}$ denote the system state and control input at each time $t\in[T]$; $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are system matrices.
@@
 Assumption 1.
-There exists $K\in\mathbb{R}^{m\times n}$ such that $A-BK$ has a spectral radius $\rho$ less than $1$.
-[^2] Thus, the Gelfand’s formula implies that there exist $C>0$, $\rho\in(0,1)$ such that $\|A-BK\|^{t}\leq C\rho^{t}$ for all $t\geq 0$.
-2.1 Model Predictive Control
+There exists $K\in\mathbb{R}^{m\times n}$ such that ABKA-BK has a spectral radius $\rho$ less than 11.222Thus, the Gelfand’s formula implies that there exist $C>0$, $\rho\in(0,1)$ such that $\|A-BK\|^{t}\leq C\rho^{t}$ for all $t\geq 0$.
+II-A Model Predictive Control
 Fix a prediction horizon to be an integer $k$ and define $\mathcal{T}\coloneqq\min\{t+k-1,T-1\}$.
@@
 $u_{t:\mathcal{T}}^{\textsc{MPC}}$ $\coloneqq\arg\min_{u}\Big{(}\sum_{\tau=t}^{\mathcal{T}}(x_{\tau}^{\top}Qx_{\tau}+u_{\tau}^{\top}Ru_{\tau})+x^{\top}_{\mathcal{T}+1}Px_{\mathcal{T}+1}\Big{)}$
-[…]
+$\text{s.t. }x_{\tau+1}=Ax_{\tau}+Bu_{\tau}+\hat{w}_{\tau|t},\forall t\leq\tau<\mathcal{T}.$ (4)
 Only the first element of $u_{t:\mathcal{T}}^{\textsc{MPC}}$ is applied as a control input to the system, and the rests are discarded.
-However, in many real-world scenarios, the disturbances are hard to predict or measure [15, 16].
+However, in many real-world scenarios, system disturbances are hard to predict [16, 17].
 In this paper, we propose a novel framework that leverages contextual information to predict future disturbances and incorporates these predictions into the decision-making process.
-2.2 Contextual MPC with Model Fine-Tuning
-[…]
+II-B Contextual MPC with Model Fine-Tuning
+Our goal is to improve the performance of the classic MPC defined in (II-A) by incorporating external contextual information at time $t\in[T]$, denoted by $c_{t}\in\mathcal{C}$ to predict future disturbances $\hat{w}_{t:\mathcal{T}|t}\in\mathcal{W}^{k}$.
+[^3] The predicted disturbance trajectory $\hat{w}_{t:\mathcal{T}|t}$ is generated in real-time, operating on the same time scale as the MPC decision-making process.
+Hence, the main challenge is to integrate a Neural Network (NN) $g_{\theta}:\mathcal{C}\to\mathcal{W}^{k}$ to detokenize the contextual information into predictions for the downstream MPC, and refine these predictions by fine-tuning the model parameter $\theta\in\Theta$ after the real disturbances are revealed.
+Let $\|\cdot\|$ denote the $\ell_{2}$-norm of a vector.
+The hypothesis set $\Theta$ has bounded diameter so that $\|\theta-\theta^{\prime}\|\leq D$ for all $\theta,\theta^{\prime}\in\Theta$ and some constant $D>0$.
+We impose the following assumptions on the NN model $g_{\theta}$, which serve as approximations for widely used architectures, including transformers [18] and input convex neural networks [19].
 Assumption 2.
@@
 $J(\theta_{1:T})-\min_{\theta\in\Theta}J(\theta),$ (5)
-where $J(\theta_{1},\ldots,\theta_{T})$ and $J(\theta^{\star})$ are the quadratic costs defined in (2a) induced by the MPC actions in (2.1) corresponding to the hyper-parameters $\theta_{1},\ldots,\theta_{T}$ and the optimal hyper-parameter $\theta^{\star}$, respectively.
-In other words, the regret in Equation 5 measures the difference between the cumulative loss incurred by $(\theta_{t}:t\in[T])$, and the cumulative loss incurred by the optimal $\theta^{\star}$ in hindsight.
+where $J(\theta_{1},\ldots,\theta_{T})$ and $J(\theta^{\star})$ are the quadratic costs defined in (2a) induced by the MPC actions in (II-A) corresponding to the hyper-parameters $\theta_{1},\ldots,\theta_{T}$ and the optimal hyper-parameter $\theta^{\star}$, respectively.
+In other words, the regret in Equation˜5 measures the difference between the cumulative loss incurred by $(\theta_{t}:t\in[T])$, and the cumulative loss incurred by the optimal $\theta^{\star}$ in hindsight.
 Next, we introduce the InstructMPC framework for contextual MPC.
-3 InstructMPC
+III InstructMPC
 In this section, we introduce InstructMPC, a framework that enables MPC controller to leverage external contextual information to effectively manage environmental disturbances.
@@
 The black lines represents the information loop, within which external contextual information is passed to the L2D module $g_{\theta}$ to produce predicted disturbances $\hat{w}_{t:\mathcal{T}|t}$.
-Then, the MPC controller $\pi_{\mathrm{MPC}}$ utilizes $\hat{w}_{t:\mathcal{T}|t}$ and the current state $x_{t}$ to determine a control input $u_{t}^{\mathrm{MPC}}$ via (2.1).
+Then, the MPC controller $\pi_{\mathrm{MPC}}$ utilizes $\hat{w}_{t:\mathcal{T}|t}$ and the current state $x_{t}$ to determine a control input $u_{t}^{\mathrm{MPC}}$ via (II-A).
 After executing the MPC control input $u_{t}^{\mathrm{MPC}}$, the environment reveals the true disturbance $w_{t}$.
 The discrepancy between $w_{t}$ and $\hat{w}_{t}$ is then sent back to the L2D module as a loss signal.
-3.1 Language to Distribution (L2D) Module
+III-A Language to Distribution (L2D) Module
 During the execution of InstructMPC, the L2D module $g_{\theta}$ predicts future environmental disturbances $\hat{w}_{t:\mathcal{T}|t}=(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})$ based on contextual information.
@@
 $\hat{w}_{t:\mathcal{T}|t}=\sum_{s\in\Omega}p(s|c_{t})w^{s}_{t:\mathcal{T}}.$
-3.2 InstructMPC Framework
+III-B InstructMPC Framework
 for $t=0,\dots,T-1$ do
  $\mathcal{T}\leftarrow\min(t+k-1,T-1)$
-       Get predictions from $g_{\theta}$ with instruction $c_{t}$: $(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})\leftarrow g_{\theta_{t}}(c_{t})$ Generate an action $u_{t}$ by solving MPC in (2.1) Update
+ Get predictions from $g_{\theta}$ with instruction $c_{t}$: $(\hat{w}_{t|t},\ldots,\hat{w}_{\mathcal{T}|t})\leftarrow g_{\theta_{t}}(c_{t})$ Generate an action $u_{t}$ by solving MPC in (II-A) Update
 $x_{t+1}$ $=Ax_{t}+Bu_{t}+w_{t}$
@@
 The L2D module is a neural network $g_{\theta}:\mathcal{C}\to\mathcal{W}^{k}$, which is updated at every time $t$ through a loss function $L_{t-k+1}:\Theta\times\mathcal{W}^{2k}\rightarrow\mathbb{R}$, which also depends on disturbances $w_{t:\mathcal{T}}$ and predictions $\hat{w}_{t:\mathcal{T}|t}$.
-An example of such a loss function can be found in Corollary 4.1 in Section 4.
+An example of such a loss function can be found in Corollary IV.1 in Section IV.
 At every time $t\in[T]$, the parameter $\theta$ of L2D module is updated according to:
@@
 In practice, when $t<k$, the NN model does not update due to insufficient observed data.
-4 Theoretical Guarantee
+IV Theoretical Guarantee
 Our main result in this section explores the online optimization of $(\theta_{t}:t\in[T])$ in the online fine-tuning process. We make the following standard assumption on the loss function $L_{t}$.
@@
 The gradient $\nabla_{\theta}L_{t}(\theta)$ is uniformly bounded, i.e., there exists $G>0$ such that by $\|\nabla_{\theta}L_{t}(\theta)\|\leq G$ for all $\theta\in\Theta$.
-Note that the loss function used for updating $g_{\theta}$ may not be aligned with regret-optimal objective that requires the knowledge of future ground-truth predictions $(w_{t}:t\in[T])$ in (2.1).
-To quantify this mismatch, we define the loss discrepancy below.
+A fundamental challenge is that the surrogate loss function used to train $g_{\theta}$ does not match the true, regret-optimal objective (II-A), which is unknown at training time.
+This misalignment creates the primary difficulty, i.e., ensuring that training the upstream L2D module $g_{\theta}$ actually improves the downstream MPC controller’s performance.
+Our solution is to define and analyze the loss discrepancy, a term that quantifies the divergence between the gradient of the L2D module surrogate loss $L$ and that of the true decision loss in Section˜II-A.
 Definition 1.
-The loss discrepancy between two loss functions $L_{1}(\theta)$ and $L_{2}(\theta)$, $\textsf{LD}(L_{1},L_{2})$ is defined as
-$\textsf{LD}(L_{1},L_{2})\coloneqq\sup_{\theta\in\Theta}\left\|\nabla_{\theta}L_{1}(\theta)-\nabla_{\theta}L_{2}(\theta)\right\|.$
-The theorem below provides a regret bound (see Equation 5) for the proposed InstructMPC, which interacts with the dynamic system described via Equation 1.
-Theorem 4.1.
+The loss discrepancy between two loss functions $L_{1}(\theta)$ and $L_{2}(\theta)$, $\emph{{LD}}(L_{1},L_{2})$ is defined as
+$\emph{{LD}}(L_{1},L_{2})\coloneqq\sup_{\theta\in\Theta}\left\|\nabla_{\theta}L_{1}(\theta)-\nabla_{\theta}L_{2}(\theta)\right\|.$
+The theorem below provides a regret bound (see Equation˜5) for the proposed InstructMPC, which interacts with the dynamic system described via Equation˜1.
+Theorem IV.1.
 Under Assumption 2,3, if the learning rate $\eta_{t}$ is non-increasing, then
 $J(\theta_{1:T})$- $J(\theta^{\star})\leq\frac{D^{2}}{2\eta_{T-1}}+\left(k-\frac{1}{2}\right)G^{2}\sum_{t=0}^{T-1}\eta_{t}$
-$+D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$ (7)
+$+D\sum_{t=0}^{T-1}\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$ (7)
 Furthermore, if we choose $\eta_{t}=\frac{D}{G\sqrt{2(2k-1)(t+1)}}$,
 $J(\theta_{1:T})$- $J(\theta^{\star})\leq 2GD\sqrt{(k-\frac{1}{2})T}$
-$+D\sum_{t=0}^{T-1}\textsf{LD}\left(L_{t},\psi_{t}^{\top}H\psi_{t}\right)+(k-1)GD,$
-where we define
+$+D\sum_{t=0}^{T-1}\emph{{LD}}\left(L_{t},\psi_{t}^{\top}H\psi_{t}\right)+(k-1)GD,$
+where $H$ is defined in Lemma 1, and we define
 $\psi_{t}(\theta)\coloneqq\sum_{\tau=t}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pg_{\theta}^{(\tau-t+1)}(c_{t}).$
 Proof.
-By Theorem 3.2 in [20], given predictions $\hat{w}_{t:\mathcal{T}|t}$, the MPC solution $u_{t}^{\textsc{MPC}}$ to the problem defined in (2.1) is
+By Theorem 3.2 in [21], given predictions $\hat{w}_{t:\mathcal{T}|t}$, the MPC solution $u_{t}^{\textsc{MPC}}$ to the problem defined in (II-A) is
 $u_{t}^{\textsc{MPC}}=-(R+B^{\top}PB)^{-1}B^{\top}\Big{(}PAx_{t}+\sum_{\tau=t}^{\mathcal{T}}(F^{\top})^{\tau-t}P\hat{w}_{\tau|t}\Big{)},$
@@
 $J(\theta_{1:T})-J(\theta^{\star})=\sum_{t=0}^{T-1}\psi_{t}(\theta_{t})^{\top}H\psi_{t}(\theta_{t})-\psi_{t}(\theta^{\star})^{\top}H\psi_{t}(\theta^{\star}),$
-where $\psi_{t}(\theta)$ is defined in Theorem 4.1.
+where $\psi_{t}(\theta)$ is defined in Theorem IV.1.
 For notational simplicity, we define $\psi_{t}\coloneqq\psi_{t}(\theta_{t})$ and $\phi_{t}\coloneqq\psi_{t}(\theta^{\star})$.
@@
 $\quad+\sum_{t=0}^{T-1}(\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})^{\top}-l_{t}^{\top})(\theta_{t}-\theta^{\star}).$ (9)
-The bound in (7) is obtained by bounding the terms in Equation 9 separately, and the details are relegated to Appendix A. ∎
-Furthermore, for a wide range of loss functions $(L_{t}:t\in[T])$, as examplified in Corollary 4.1 below, they exhibit a bounded discrepancy from the MPC cost with the discrepancy decaying exponentially as $k$ increases, i.e., $\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k}$.
-Corollary 4.1.
+The bound in (7) is obtained by bounding the terms in Equation˜9 separately, and the details are relegated to Appendix Appendix A: Proof of Theorem IV.1. ∎
+Furthermore, for a wide range of loss functions $(L_{t}:t\in[T])$, as examplified in Corollary IV.1 below, they exhibit a bounded discrepancy from the MPC cost with the discrepancy decaying exponentially as $k$ increases, i.e., $\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k}$.
+Corollary IV.1.
 Under Assumption 1, for the case when $L_{t}(\theta)=\hat{\psi}_{t}(\theta)^{\top}H\hat{\psi}_{t}(\theta)$ with
 $\hat{\psi}_{t}(\theta)\coloneqq\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\sum_{\tau=t}^{\mathcal{T}}\left(F^{\top}\right)^{\tau-t}Pg_{\theta}^{(\tau-t+1)}(c_{t}),$
-there exist constants $C$ and $10<\rho<$ 1 such that for all $t\in[T]$,
-$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k},$
-where $\psi_{t}(\theta)$ is defined in Theorem 4.1.
-For the particular choice of $L_{t}$ in Corollary 4.1, the term $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})$ in the bound (7) of Theorem 4.1 simplifies to $CDT\rho^{k}$.
+there exist constants $C$ and $0<\rho<1$ such that
+$\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq C\rho^{k},\ \text{ for all }t\in[T],$
+where $\psi_{t}(\theta)$ is defined in Theorem IV.1.
+For the particular choice of $L_{t}$ in Corollary IV.1, the term $D\sum_{t=0}^{T-1}{\texttt{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})$ in the bound (7) of Theorem IV.1 simplifies to $CDT\rho^{k}$.
 This provides an insightful guideline for selecting $k$.
 Specifically, by setting $k=\log_{1/\rho}T,$ we can achieve a regret bound of $O(\sqrt{T\log T})$.
-5 Numerical Experiment
-In this section, we present a detailed demonstration of how the L2D module is fine-tuned in Subsection 5.1.
+V Numerical Experiment
+In this section, we present a detailed demonstration of how the L2D module is fine-tuned in Subsection V-A.
 We implement the L2D module with LLaMa-8B [22], a causal language model.
@@
 To obtain these probabilities, we use the following prompt template: “[task description].
-Given the current context: [context description], the expected scenario is [MASK]." For a predefined set of scenarios ($\mathcal{S}$ in Subsection 3.1), we evaluate the prediction score assigned by the model by inserting each candidate token into the [MASK] position.
+Given the current context: [context description], the expected scenario is [MASK].” For a predefined set of scenarios ($\mathcal{S}$ in Subsection III-A), we evaluate the prediction score assigned by the model by inserting each candidate token into the [MASK] position.
 These scores are then normalized using a softmax function to produce a probability distribution.
-In Subsection 5.2 and Subsection 5.3, we demonstrate the effectiveness of InstructMPC framework in two scenarios.
+In Subsection V-B and Subsection V-C, we demonstrate the effectiveness of InstructMPC framework in two scenarios.
 The first is a robot tracking task, where an agent is required to follow an unknown trajectory.
@@
 The second scenario is a real-world energy management task, in which human-provided instructions, such as news and scheduled appointments, are used to predict external disturbances.
-In this case, we further extend InstructMPC to operate on a nonlinear dynamical system, we also adopt a piecewise linear cost instead of the quadratic cost defined in eq.
-2a, demonstrating the generalizability of InstructMPC in diverse system dynamics and different cost functions.
-5.1 Fine-Tuning L2D Module
+In this case, we further extend InstructMPC to operate on a nonlinear dynamical system, we also adopt a piecewise linear cost instead of the quadratic cost defined in eq.˜2a, demonstrating the generalizability of InstructMPC in diverse system dynamics and different cost functions.
+V-A Fine-Tuning L2D Module
 Fine-Tuning L2D Module.
@@
 Specifically, the parameter $\theta$ is updated when $|\mathcal{D}|=240$ in the robot tracking task and $|\mathcal{D}|=500$ in the energy management scenario.
-5.2 Application 1: Robot Tracking
+V-B Application 1: Robot Tracking
 Problem description. We examine a 2D robotic tracking scenario [23, 21] where a controller follows an unknown, fixed astroid path, which is
@@
 $A\coloneqq\begin{bmatrix}1&0&0.2&0\\ 0&1&0&0.2\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix},B\coloneqq\begin{bmatrix}0&0\\ 0&0\\ 0.2&0\\ 0&0.2\end{bmatrix}$,
-$w_{t}\coloneqq Ay_{t}-y_{t+1}+Z_{t},$
-$Z_{t}\coloneqq\begin{bmatrix}0\!\!\!&0\!\!\!&-0.2Z_{t}^{(1)}\!\!\!&-0.2Z_{t}^{(2)}\end{bmatrix}^{\top}\!\!\!,$
+$w_{t}\coloneqq Ay_{t}-y_{t+1}+Z_{t},Z_{t}\coloneqq\begin{bmatrix}0\!\!\!&0\!\!\!&-0.2Z_{t}^{(1)}\!\!\!&-0.2Z_{t}^{(2)}\end{bmatrix}^{\top}\!\!\!$,
 where $Z_{t}$ is a random variable representing the random wind force.
@@
 $\begin{cases}Z^{(1)}_{t},Z^{(2)}_{t}\sim\mathcal{U}(-45,45)&t\in T_{\text{sub}}\\ Z^{(1)}_{t},Z^{(2)}_{t}\sim\mathcal{U}(-2,2)&\text{otherwise}\end{cases}$
-To track the trajectory, the controller uses quadratic cost in Equation 2a with
+To track the trajectory, the controller uses quadratic cost in Equation˜2a with
 $Q\coloneqq\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{bmatrix},\ \ R\coloneqq\begin{bmatrix}10^{-2}&0\\ 0&10^{-2}\end{bmatrix}$.
@@
   1. 1.
-    Classic MPC: An MPC baseline using (2.1) where disturbance predictions are set to zero ($\hat{w}_{t:\mathcal{T}|t}=0$) for all $t$, equivalent to a Linear Quadratic Regulator (LQR) for the problem (2a)-(2b).
+    Classic MPC: An MPC baseline using (II-A) where disturbance predictions are set to zero ($\hat{w}_{t:\mathcal{T}|t}=0$) for all $t$, equivalent to a Linear Quadratic Regulator (LQR) for the problem (2a)-(2b).
   2. 2.
@@
 In contrast, both methods with L2D predictions improve the performance of the robot, with the fine-tuned version demonstrating a closer robot path to the designed trajectory.
-Figure 2: Application 1: Robot Tracking (Section 5.2).
+Figure 2: Application 1: Robot Tracking (Section˜V-B).
 Average robot tracking trajectories of three methods: 1.
-MPC in (2.1) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2.
+MPC in (II-A) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2.
 MPC with predictions from a zero-shot prompting LLaMA-8B [22] model; 3.
 InstructMPC with LLaMA-8B [22] fine-tuned in a closed-loop.
-For each method, we ran $20$ episodes, and plot the average robot trajectory.
+For each method, we ran 2020 episodes, and plot the average robot trajectory.
 The shadow region denotes the variance of trajectories.
-5.3 Application 2: Building Energy Management
+V-C Application 2: Building Energy Management
 Problem description.
@@
 These disturbances are inferred from contextual information, including weather forecasts, human activity schedules, and natural language instructions.
-Examples of such contextual information include: “The CDC conference deadline is approaching, and many people are running experiments," or “I will be fine-tuning my LLaMA-1B model at 3 PM today, which will take approximately two hours." The objective of the energy management system is to minimize the total cost of purchasing electricity from the grid.
+Examples of such contextual information include: “The CDC conference deadline is approaching, and many people are running experiments,” or “I will be fine-tuning my LLaMA-1B model at 3 PM today, which will take approximately two hours.” The objective of the energy management system is to minimize the total cost of purchasing electricity from the grid.
 This cost is computed using a piecewise linear function that reflects real-world peak and valley electricity prices.
-Figure 3: Application 2: Building Energy Management (Section 5.3).
+Figure 3: Application 2: Building Energy Management (Section˜V-C).
 Daily electricity cost comparison of three approaches: 1.
-MPC in (2.1) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2.
+MPC in (II-A) with predictions $\hat{w}_{t:\mathcal{T}|t}=0$ for all $t$ (classic MPC); 2.
 MPC with predictions from a zero-shot prompting LLaMA-8B [22] model; 3.
@@
 The upper-right part provides a summary table comparing the total electricity cost and the corresponding percentage reduction, while the bottom-right part provides a closer view of daily costs over $40$ days, further illustrating the cost savings by InstructMPC.
-6 Conclusion
+VI Conclusion
 We have presented the InstructMPC framework that integrates real-time human instructions into Model Predictive Control (MPC) via a human-LLM-in-the-loop approach.
@@
 For linear dynamics, we established a robust performance guarantee, proving that IMPC achieves a regret bound of $O(\sqrt{T\log T})$.
-In future work, we aim to extend these results to nonlinear systems and further explore the integration of richer contextual cues to enhance adaptability and robustness in complex environments.
-Appendix A Proof of Theorem 4.1
-The following lemma characterizing the quadratic cost gap is useful.
+In future work, we aim to extend these results to nonlinear systems, and plan to conduct more experiments under different key parameters, such as the prediction horizon $k$ and the choice of LLM.
+We will also deploy the InstructMPC framework on our newly deployed physical energy testbed OpenCEM [25] to assess real-world generalization beyond synthetic workloads.
+Appendix A: Proof of Theorem IV.1
+The following lemma characterizes the quadratic cost gap.
 Lemma 1 (Lemma 13 in [26]).
-For any $\psi_{t}\in\mathbb{R}^{n}$, if at each time $t\in[T]$, a controller $\pi$ implements a control input
-$u_{t}^{\pi}=-(R+B^{\top}PB)^{-1}B^{\top}\Big{(}PAx_{t}+\sum_{\tau=t}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\psi_{t}\Big{)},$
-then the difference between the optimal cost $J^{\star}$ and the algorithm cost $J(\pi)$ is given by:
-$J(\pi)-J^{\star}=\sum_{t=0}^{T-1}\psi_{t}^{\top}H\psi_{t},$
-where $H\coloneqq B(R+B^{\top}PB)^{-1}B^{\top}$ and $F\coloneqq A-HPA$.
-Now, we bound the terms in Equation 9 separately:
+For any $\psi_{t}\in\mathbb{R}^{n}$, if at each time $t\in[T]$, a controller $\pi$ implements a control input $u_{t}^{\pi}=-(R+B^{\top}PB)^{-1}B^{\top}\big{(}PAx_{t}+\sum_{\tau=t}^{T-1}\left(F^{\top}\right)^{\tau-t}Pw_{\tau}-\psi_{t}\big{)},$ then the difference between the optimal cost $J^{\star}$ and the algorithm cost $J(\pi)$ is given by $J(\pi)-J^{\star}=\sum_{t=0}^{T-1}\psi_{t}^{\top}H\psi_{t},$ where $H\coloneqq B(R+B^{\top}PB)^{-1}B^{\top}$ and $F\coloneqq A-HPA$.
+Now, we bound the terms in Equation˜9 separately:
 $\sum_{t=0}^{T-1}\nabla_{\theta_{t}}(\psi_{t}^{\top}H\psi_{t})^{\top}(\theta_{t}-\theta^{\star})$
@@
 $=$ $\|\theta_{t}-\theta^{\star}\|^{2}-2\eta_{t}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})+\eta_{t}^{2}\|l_{t-k+1}\|^{2}.$ (10)
-Then, rearranging the terms in Equation 10 and noting that the gradients are bounded by Assumption 3,
-$2l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})\leq\frac{\|\theta_{t}-\theta^{\star}\|^{2}-\|\theta_{t+1}-\theta^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (11)
-Summing above from $t=0$ to $T-1$, for (a),
-22 22 $\sum_{t=0}^{T-1}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})$
+Then, rearranging the terms in Equation˜10 and noting that the gradients are bounded by Assumption 3,
+$2l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})\leq\frac{\|\theta_{t}-\theta^{\star}\|^{2}-\|\theta_{t+1}-\theta^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$
+Summing above from $t=0$ to T1T-1, for (a),
+$2$ $\sum_{t=0}^{T-1}l_{t-k+1}^{\top}(\theta_{t}-\theta^{\star})$
 $\leq$ $\sum_{t=0}^{T-1}\frac{\|\theta_{t}-\theta^{\star}\|^{2}-\|\theta_{t+1}-\theta^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}$
@@
 $\leq$ $\frac{D^{2}}{\eta_{T-1}}+G^{2}\sum_{t=0}^{T-1}\eta_{t}.$ (11)
-Furthermore, the second term in Equation 9 can be bounded from above as
-[…]
+Furthermore, the second term in Equation˜9 can be bounded from above as
+$\text{(b)}=$ $\sum_{t=0}^{T-1}(l_{t}-l_{t-k+1})^{\top}(\theta_{t}-\theta^{\star})$
 $=$ $\sum_{t=0}^{T-k}l_{t}^{\top}(\theta_{t}-\theta_{t+k-1})+\sum_{t=T-k+1}^{T-1}l_{t}^{\top}(\theta_{t}-\theta^{\star})$
@@
 $\leq$ $\frac{D^{2}}{2\eta_{T-1}}+(k-\frac{1}{2})G^{2}\sum_{t=0}^{T-1}\eta_{t}+(k-1)GD$
-$+$ $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t}),$ (15)
-where we have used (12) and (13) to derive  (14); the last inequality (15) holds by the loss discrepancy in Definition 1. Finally, applying $\eta_{t}={D}/\left({G\sqrt{2(2k-1)(t+1)}}\right)$ to (15),
+$+$ $D\sum_{t=0}^{T-1}\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t}),$ (14)
+where we have used (11) and (12) to derive  (13); the last inequality (14) holds by the loss discrepancy in Definition˜1. Finally, applying $\eta_{t}={D}/\left({G\sqrt{2(2k-1)(t+1)}}\right)$ to (14),
 $J(\theta_{1:T})-J(\theta^{\star})$
 $=$ $GD\sqrt{(k-\frac{1}{2})T}+\sqrt{k-\frac{1}{2}}\frac{GD}{2}\sum_{t=0}^{T-1}\frac{1}{\sqrt{t+1}}$
-$+$ $D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD$
-$\leq$ $2GD\sqrt{(k-\frac{1}{2})T}+D\sum_{t=0}^{T-1}\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$ (16)
-Inequality (16) in the second statement of Theorem 4.1 is obtained by using $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$.
-Appendix B Proof of Corollary 4.1
+$+$ $D\sum_{t=0}^{T-1}\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD$
+$\leq$ $2GD\sqrt{(k-\frac{1}{2})T}+D\sum_{t=0}^{T-1}\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})+(k-1)GD.$
+The inequality above in the second statement of Theorem IV.1 is obtained by using $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\leq 2\sqrt{T}$.
+Appendix B: Proof of Corollary IV.1
 Suppose $L_{t}(\theta)=\hat{\psi}_{t}(\theta)^{\top}H\hat{\psi}_{t}(\theta)$ where
@@
 By Definition 1,
-$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})$ $=\left\|\frac{\partial(\psi_{t}^{\top}H\psi_{t})}{\partial\theta}-\frac{\partial(\hat{\psi}_{t}^{\top}H\hat{\psi}_{t})}{\partial\theta}\right\|$
+$\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})$ $=\left\|\frac{\partial(\psi_{t}^{\top}H\psi_{t})}{\partial\theta}-\frac{\partial(\hat{\psi}_{t}^{\top}H\hat{\psi}_{t})}{\partial\theta}\right\|$
 $=\left\|2(\psi_{t}-\hat{\psi}_{t})^{\top}H\frac{\partial\psi_{t}}{\partial\theta}\right\|$ (15)
@@
 We have use the fact that $\frac{\partial\psi_{t}}{\partial\theta}=\frac{\partial\hat{\psi}_{t}}{\partial\theta}$ to get (15).
-Note that if $\mathcal{T}=T-1$, then $\psi_{t}=\hat{\psi}_{t}$, $\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})=0$.
+Note that if $\mathcal{T}=T-1$, then $\psi_{t}=\hat{\psi}_{t}$, $\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})=0$.
 Thus, it suffices to restrict our analysis to the case when $\mathcal{T}=t+k-1$.
@@
 $\left\|\frac{\partial\psi_{t}}{\partial\theta}\right\|$ $=\left\|\sum_{\tau=t}^{t+k-1}\left(F^{\top}\right)^{\tau-t}P\nabla_{\theta}g_{\theta_{t}}^{(\tau-t+1)}(c_{t})\right\|$
-$\leq L\|P\|\sum_{\tau=t}^{t+k-1}\|F\|^{\tau-t}$
-$\leq L\frac{\|P\|}{1-\|F\|}.$ (21)
-Finally, combining (18),(20), and (21) we have
-$\textsf{LD}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq 2\frac{LW\|P\|^{2}\|H\|\|F\|^{k}}{(1-\|F\|)^{2}}.$ (22)
+$\leq L\|P\|\sum_{\tau=t}^{t+k-1}\|F\|^{\tau-t}\leq L\frac{\|P\|}{1-\|F\|}.$
+Finally, combining (16),(18), and (Appendix B: Proof of Corollary IV.1) we have
+$\emph{{LD}}(L_{t},\psi_{t}^{\top}H\psi_{t})\leq 2\frac{LW\|P\|^{2}\|H\|\|F\|^{k}}{(1-\|F\|)^{2}}.$ (19)